OpenStudy (anonymous):

i forgot...how would you simplify 64^-1/3

OpenStudy (john_es):

\[64^{-1/3}=(4^3)^{-1/3}=(4)^{-3/3}=4^{-1}\]

OpenStudy (anonymous):

it looks like ok thank you